FEL Upgrade Project Weekly Brief
October 29-November 2, 2001



Highlights:
The 3rd 100 kW was received from the vendor this week, which completes this contract.  The 1st 100 kW
klystron and solenoid were installed on their cart and installation of the LCW was started.  Vacuum of all three
(3) klystrons proved to be good.

Management:
Preparations continued for the November 8-9 semiannual review of the FEL Upgrade project.

WBS 4 (Injector):
Gun HVPS - The design of the HVPS Tank is making good progress.  A design for a combined conditioning &
running resistor is in progress.  A new combined resistor would eliminate the transmission line elbow and
hangers.

WBS 6 (RF):
Zone 3 - Essentially complete.  Missing a key lock on the HVPS and minor testing of tuners.

Zone 4 - Essentially complete.  Missing a panel, klystron and circulator final testing of position 8, and the
correct PS for the 5 watt amplifiers.

Quarter - All 3 of the 100 kW klystrons have been received from CPI.  The first one is installed with its
solenoid on the klystron cart.  To avoid a sharp bend, the LCW piping is being modified slightly.  The design of
the HVPS continues.

WBS 8 (Instrumentation):
Two man-days of operations support in the form of duty officers was provided during the week.

The Laser web site upgrades continue.  A new format for the home page will include photos from the Photo
Gallery at log on.  The task list continues to become more functional and minor improvements are making it
more user friendly.

New stepper motors have been received and modified to work with existing chassis.  A new control circuit for
these motors has been developed and is being tested in parallel.

Development and evaluation of a vacuum control system for SNS continues.  The first of the new Ion Pump
Power Supplies have been received and are being tested and evaluated.

The design phase of the Digital Isolation Module and the TTL Fan-Out Module is complete.  The TTL Fan-Out
Module drawing package is also complete and these boards are ready to be sent out for production.  Minor
modifications have been made to the General Purpose Carrier Board design and the Programmable Time and
Logic Board design.  EECAD is modifying the drawings to include the changes for the next production cycle.

Welcome to our newest I & C Team member, Jim Coleman who comes to us from the safety group.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
  We plan to inspect the first coil potting in two weeks at Magnet Enterprises International.
Injector Dipoles (DU/DV)
  DULY continues to revise small portions of the final assemblies that need modification.
Arc Dipoles (GY, GX, GQ)
  Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ)
    and bringing the faces and the gap in conformance with the magnetic model.
The 180° Dipole (DY) is in its procurement cycle
Magnet Measurement
  The drawings for the measurement stands for the two quadrupoles are being made.
  The parts for the dipole measurement stand are in detailing.
  We continued work on the probe that will measure the sextupole and the octupole.
Quadrupoles
3 inch quad (QX)
  The sample aluminum version of the core is machined and is being inspected before finish machining of the
    first article will be complete.
  The assembly vendor fed back to us that the coils were not potted in the style he was anticipating.  After
    some exchanging of photographs and discussions with their manufacturer, Magnet Enterprises International,
    the question was resolved.  However, we are sending Kevin Sullivan to inspect the coils and consult with New
    England Techni Coil about the details of assembly.
Trim Quad (QT)
  After the discussions about our QX coils, Tom Hiatt determined items he wanted to change in the
    manufacturing specification.  After the revision, he will go out with the final bid cycle.
Sextupole (SF)
  We are composing the statement of work to DULY Research to model the sextupole.
Octupole (OT)
  Work on this magnet is on furlough.
Corrector Dipoles (DB, DJ)
  Milhous Co. delivered another batch of DB coils and qualified the DJ first article.
Beam Line and Vacuum
  Dave Waldman completed the task order for design of the Arc Chambers utilizing the Engineering Services
    contract with Advanced Energy Systems.
  Design work on girders and stands continued at a high level on the quadrupole telescope in front of the first
    arc and the region between the cryomodules.

WBS 10 (Wiggler):
Detailed measurements of the wigglers continue, along with the analysis of the extensive data set taken to
date.  So far, the wiggler quality looks excellent with the wiggler exceeding its specifications nicely.

Tommy Hiatt is trying to get the lift hoists for the wiggler and dispersion section fully qualified by safety
personnel so that we can fit up the vacuum chamber next week.

WBS 11 (Optics):
We are continuing to detail components of the optical cavity, with emphasis currently on the mirror mounting
and the yaw yoke "big C".  More components (flexural pivots and couplings) for the optical cavity have
arrived.  We now have design support for the beamline components for the active mirror steering system.
Assembly of the ultraviewer began this week with component cleaning.  Preparations continue for the
upcoming semiannual review on Nov. 8.
 
The following work was done to support operations:

We provided support for our FEL experiments, specifically, setting up optics to attempt to measure the beam
quality under macropulsed operation, and setting up equipment for the upcoming THz experiments.  As an
update on last week's attempt to find the laser-induced damage to the 3 um coatings, we were unable to create
damage.

We welcome Dr. Shukui Zhang to the FEL Dept.  He will be working in the Optics Group on new diagnostics
for the FEL.

Michelle Shinn presented a paper on the FEL upgrade path to participants of the 4th annual Directed Energy
Professionals Society conference in Huntsville, AL.

Operations/Commissioning:
This past week four (4) different users were provided beam at two (2) different wavelengths where they
successfully ran their experiments.  M. Kelley ran his Rapid Thermal Processing of Polymer experiment.  S.
Benson ran FEL beam tests with ZnSe mirrors.  L. Merminga et al. ran Beam Break Up tests for two (2) days
and will continue this weekend with HOM Load tests.  M. Shinn ran her Absorption of FEL Mirrors at 5 um
experiment. The results of the last 3 experiments will be factored into the design of the upgrade machine.